Consider the following information:

  

Rate of Return if State Occurs
  State of Probability of
  Economy State of Economy Stock A Stock B Stock C
  Boom .20 .38 .48 .28
  Good .50 .14 .19 .12
  Poor .20 .05 .08 .06
  Bust .10 .19 .23 .09

  

a.

Your portfolio is invested 22 percent each in A and C, and 56 percent in B. What is the expected return of the portfolio? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  Expected return %

  

b-1. What is the variance of this portfolio? (Do not round intermediate calculations and round your answer to 5 decimal places, e.g., 32.16161.)
  Variance   
b-2.

What is the standard deviation? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  Standard deviation %
Answer #1

State Portfolio Return
0.20 =22%*0.38+56%*0.48+22%*0.28=41.40000%

0.50 =22%*0.14+56%*0.19+22%*0.12=16.36000%

0.20 =22%*(-0.05)+56%*(-0.08)+22%*(-0.06)=-6.90000%

0.10 =22%*(-0.19)+56%*(-0.23)+22%*(-0.09)=-19.04000%

1.
=0.20*41.40%+0.50*16.36%+0.20*(-6.90%)+0.10*(-19.04%)
=13.17600%

2.
=0.20*(41.40%-13.17600%)^2+0.50*(16.36%-13.17600%)^2+0.20*(-6.90%-13.17600%)^2+0.10*(-19.04%-13.17600%)^2
=0.034878398

3.
=sqrt(0.034878398)
=18.67576%